Heavy Products, Inc. developed standard costs for direct material and direct labor. In 2017, AII estimated the following standard costs for one of their major products, the 10-gallon plastic container.

                                                Budgeted quantity           Budgeted price

        Direct materials         0.30 pounds                        $90 per pound

        Direct labor                 0.25 hours                            $30 per hour

During June, Heavy Products produced and sold 15,000 containers using 2,400 pounds of direct materials at an average cost per pound of $92 and 3,750 direct manufacturing labor-hours at an average wage of $91.25 per hour.

The direct manufacturing labor efficiency variance during June is ________.

Group of answer choices

$918,750 unfavorable

$229,687.5 favorable

$0

$85,547 unfavorable
Answer #1

Labor efficiency variance

= (Standard hours - Actual hours) * standard rate

= (15,000*0.25 - 3,750) * 30

= (3,750 - 3,750) * 30

= 0 * 30

= 0